OpenStudy (migitmack):

Your daily workout plan involves a total of 60 minutes of jogging and biking. You burn 10 calories per minute jogging and 5 calories per minute biking. How many calories will you burn in your 60−minute workout if you jog for 30 minutes?

OpenStudy (btaylor):

30(10) + 30(5) = 300+150 = 450 calories.